OpenStudy (anonymous):

The sum of two numbers is 52 . One number is 3times as large as the other. What are the numbers?

OpenStudy (anonymous):

x + y = 52 x = 3y substitute the second equation into the first one and solve

OpenStudy (anonymous):

what number did you get that went into 52 3times

OpenStudy (anonymous):

i didn't do that. i plugged the second equation in the first one. so x + y = 52 becomes: 3y + y = 52 4y = 52 y = 13 plug y in the second equation (which is x = 3y) x = 3(13) = 39